A car travels 40 kilometers at an average speed of 80 kilometers per hour and then travels 40 kilometers at an average speed of 40 kilometers per hour. The average speed of the car for this 80-kilometer trip is?
1
Expert's answer
2019-10-07T10:59:17-0400

"\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ t=t_1+t_2"

Where,

"\\t_1=time\ elapsed \ for first\ half""t_2=time \ for\ second\ half"

t = time with average speed

"\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \frac{80}{v_{avg}}=\frac{40}{80}+\frac{40}{40}=\frac{3}{2}\\\ \ \ \ \ \ \ \ \ \ \ \ \ \ v_{avg}=\frac{2\times 80}{3}=53.33\ m\ s^{-1}"